OpenStudy (anonymous):

Find the number of permutations of the first ten letters of the alphabet taking two letters at a time

OpenStudy (nurali):

\[nPr=\frac{ n! }{(n-r)! }\] \[n=10\] and \[r=2\] \[10P2=\frac{ 10! }{ (10-2)! }\] \[10P2=\frac{ 10*9*8! }{ 8! }\] \[10P2=10*9\] \[10P2=90\] 90 Ways